OpenStudy (anonymous):

Solve the proportion 2 times x all over 9 times x minus 8 equals 2 over x.. What is the value of x? A:x = 1 and x = 9 B:x = −1 and x = −9 C:x = 1 and x = 8 D:x = −1 and x = 8

OpenStudy (tkhunny):

\(\dfrac{2x}{9x-8} = \dfrac{2}{x}\) Is that what you intended? What's your best work toward solving it?

OpenStudy (anonymous):

Nop it is i yhink it is 9(x-8)

OpenStudy (anonymous):

@UditKulka :Yes ! I think so ! we have : 2x/9(x-8)=2/x It s our equation . K ? @tswar: ??

OpenStudy (anonymous):

The question is unclear that is what i can say

OpenStudy (tkhunny):

Write down \(x\ne 0\) and then multiply by x.

OpenStudy (anonymous):

I'm really confused now...